# DP-Fusion: Token-Level Differentially Private Inference for Large Language Models

### v3

@@ 4.3 Empirical Privacy Attacks
Section˜4.2 gives an upper bound on the attacker’s advantage measured in the token recovery game, whereas this section describes empirical attacks to measure lower bounds. The token-recovery game states that when attacking a privacy group j{j}, the attacker’s goal is to predict, from the candidate set $C_{j^{*}}$, which (ordered) token set was present in the original input document $D$ which was used as an input to produce the privatized document $D^{\prime}$. This is analogous to prior Membership Inference Attacks (MIA) on LLMs and we can evaluate prior works such as the Min-K Attack (shi2023detecting), and the standard baseline LOSS Attack (yeom2018privacy), described as follows.
Min-K Attack (shi2023detecting). The inference attack, Min-K% shi2023detecting, calculates the average log-likelihood of the k% least-probable tokens i.e., the tokens with the lowest predicted probabilities $\Pr(x_{i}\mid x_{<i})$ in a sequence $x=(x_{1},x_{2},\dots,x_{N})$: $\text{MIN-K\% PROB}(x)=\frac{1}{|\text{Min-K\%}(x)|}\sum_{x_{i}\in\text{Min-K\%}(x)}\log\,\Pr\bigl(x_{i}\,\big|\;x_{1},\dots,x_{i-1}\bigr).$
The LOSS attack (yeom2018privacy). This attack calculates a loss for each of the candidate secrets using a surrogate model and the paraphrased document and uses it as a score to predict the true secret.
@@ 5 Experiments
Our experiments use the Qwen 2.5 7B-Instruct model (qwen2025qwen25technicalreport) running on a single A100 GPU. We replicate the DPI baseline methods DP-Decoding and DP-Prompt using their publicly released code. To provide a comprehensive overview, we measure utility with multiple metrics: (i) perplexity, computed via teacher forcing on the ground truth document $D$, and (ii) LLM-as-a-judge win-rate, with GPT-4o-mini judge, to compare pairs of generated paraphrases from different methods. We measure privacy through: (i) an upper bound with theoretical guarantees $(\epsilon)$ and (ii) as a lower bound through the adversary’s success rate in the token-recovery game.
@@ 5.1 Experimental Setup
We focus on TAB-ECHR (pilan2022text) which is a hand-annotated collection of European Court of Human Rights (ECHR) cases (echr_dataset), where private information of eight types (PERSON, CODE, LOC, ORG, DEM, DATETIME, QUANTITY, MISC) is marked. We refer to Section˜A.4 in the Appendix for more details.
@@
2) Empirical Defenses. To simulate simple NER and prompt-engineering baselines (Sec. 2.1), we include two other defenses: No DPI - NER and No DPI - Original Document, where the LLM directly paraphrases the document using only the public tokens $X_{p}$ or the full prompt $D$, respectively (Sec. 4.1). As such approaches will typically involve manually updating the prompt to improve privacy in practical settings, we modify the base prompt (Appendix A.7) with instructions like “produce a natural paraphrase of this for ensuring privacy.” For TAB-ECHR (Sec. 5.1), private tokens are already hand-labeled, so we use these labels directly instead of running an NER system.
@@ 5.3 Utility measured by LLM-as-a-Judge
While perplexity measures token-level fit on the original document $D$, it does not reflect the quality of the generated paraphrase $D^{\prime}$ (Examples in Appendix A.8). Hence, we also evaluate utility with the LLM-as-a-judge setup gu2025surveyllmasajudge We provide the judge, GPT-4o-mini, with the original document and a pair of paraphrases from different methods or settings, and prompt it (Appendix A.9) to select the paraphrase that retains more information from the original document. We report the resulting win rates in Figure 4, with full support counts for the comparisons in Appendix A.10. For DP-Prompt, we only report the results with $width=50$, as the $width=5$ setting consistently yields garbled outputs, both upon inspection (Appendix A.8) and as indicated by the high perplexity score (Figure 5). DP-Fusion substantially improves over other DPI baselines in this evaluation. Even at the strong privacy (lowest utility setting) ($\alpha\beta=0.01$), it outperforms DP-Decoding and DP-Prompt with $\geq 95\%$ win rate on all settings, except DP-Prompt at $T=0.75$. However, this setting of DP-Prompt is unusable in privacy-focused scenarios, as it provides very low empirical privacy, which we observe in the following section. DP-Fusion, surpasses the public baseline $45\%$ of the time, and at $\alpha\beta=0.1$, it exceeds the public baseline (56% win rate). Within DP-Fusion, stronger privacy ($\alpha\beta=0.01$) yields a lower LLM-as-a-judge win rate than weaker privacy ($\alpha\beta=0.1$), illustrating the expected privacy/utility trade-off.
@@ 5.4 Lower Bounds on Privacy
The ASR of the attacks described in Sec. 4.3, together with the corresponding perplexity values of each method (Sec. 5.2), are showcased in Table 1. For each defense method, we report results in two configurations: the highest-utility (lowest-privacy) and the lowest-utility (highest-privacy) settings, as implemented in Section 5.1. We can see that DP-Fusion achieves a $6\times$ higher utility as the best baseline DPI method DP-Prompt at comparable privacy levels (1.426 for $\alpha\beta_{i}=0.10$ versus 8.44 for $\text{w=}50,\text{T=}1.75$). Full results across all parameter settings are presented in Appendices A.11, A.13, and A.12. Additionally, ASR versus $\epsilon$ plots are in Appendix A.14 and A.15.
@@
LOSS-based attack has the highest ASR across all settings. In the strictest privacy setting ($\alpha\beta_{i}=0.01$), DP-Fusion achieves a perplexity (1.459), which is nearly identical to the No DPI - NER baseline (1.46), while maintaining a lower ASR (0.26 vs. 0.2767), thereby offering slightly better utility/privacy tradeoff, but with formal DP guarantees. In the more relaxed setting ($\alpha\beta_{i}=0.10$), DP-Fusion improves utility (PPL = 1.426) with only a marginal increase in ASR (+3.3%). On the other-hand, baseline DPI methods, DP-Decoding and DP-Prompt exhibit significantly higher perplexity (e.g., >100 for DP-Prompt with width 5), indicating heavily degraded outputs. Although DP-Prompt with width 50 and $T=0.75$ achieves lower perplexity (4.26) and produces good quality paraphrases (Figure 4), it does so at the cost of high $\varepsilon$ values (>100,000) and ASR (around 50%), thus providing almost no formal or empirical privacy guarantee.
@@ 6 Discussion
Role of Tagging. DP-Fusion assumes a tagger to define privacy groups; its DP guarantees apply to the tagged spans. Thus, low false negatives (FN) are required for coverage, not for the validity of the mechanism. On TAB-ECHR, off-the-shelf taggers (microsoft_presidio_docs), already achieve low FN ($3.9\%$ FN, $85.4\%$ F1, Appendix A.16) and can be tuned further. With real-world taggers in pipeline, the gap between DP-Fusion and No-DPI NER widens, since we can compensate for tagger imperfections by tuning assigned $\epsilon$ (Appendix A.17). Therefore, developing PII taggers is orthogonal to our work, and DP-Fusion benefits from developments in better NER systems.
@@
Single-Group Implementation. Although we support per-group $\epsilon$, this requires computing m+1m{+}1 distributions per step (1 public + $m$ private), making inference $\approx(m{+}1)\times$ heavier in memory and compute. Increasing $m$ tightens the theoretic privacy (per-group $\epsilon$ decreases with $m$; Thm. 4), but it also increases the effective weight of the public distribution in $p_{\text{final}}$, i.e., more of the public view leaks through. In practice, these effects make the multi-group variant less smooth: as $m$ grows, the fused distribution is increasingly dominated by $p_{\text{pub}}$, so the transition from paraphrasing $D$ to paraphrasing $D\setminus T_{\text{priv}}$ does not vary smoothly with $\epsilon$. We therefore also implement single-group DP-Fusion with one shared $\epsilon$ (Appendix A.18). This variant is more efficient and yields a smoother privacy–utility curve at the expense of weaker theoretical guarantees. We hypothesize that dataset characteristics also contribute. On a different medical PII dataset with more private tokens that impact output paraphrases maccrobat2020, we observe smoother privacy–utility trade-offs and a larger gap to No-DPI NER baseline (Appendix A.19).
@@
DP-Fusion against Prompt Injection Attacks. LLMs can be augmented by external databases (e.g., via RAG or web search tools). Given a query, they retrieve multiple chunks from different, potentially untrustworthy sources, which makes the LLM vulnerable to prompt injection attacks (liu2024formalizing). We use a RAG pipeline and poison a single chunk to jailbreak, achieving an attack success rate (ASR) of $\geq 90\%$ against undefended models. Since the provider knows which chunks came from which source, they can label each chunk as a ’privacy group’ and provably bound the influence of any chunk using DP-Fusion. The defender can control a security/utility trade-off against prompt injection that gracefully degrades toward the no-defense level for larger $\alpha\beta$ values. We evaluate DP-Fusion for different $\alpha\beta$ values and find that at 0.001, 0.01, and 1.0 it provides perfect security ($0\%$ ASR). Full details are in Appendix A.20.
@@ Appendix A Appendix
A.1 Bisection Search
@@
A.2 Proof of Monotonicity of Rényi Divergence
@@
A.3 Monotonicity of Divergence in $\lambda$
@@
A.4 Detailed information about the TAB-ECHR dataset
@@
  • Quasi identifiers: Publicly known information that doesn’t enable re-identification in isolation but may do so when combined with other quasi-identifiers in the same context. For example, the combination of gender, birth date, and postal code can uniquely identify 63-87% of the U.S. population golle2006revisiting.
@@
A.5 Ablation on candidate set size
@@
A.6 Comparison between data dependent and theoretical $\epsilon$, from $\alpha\beta$
@@
A.7 Paraphrasing Prompt Template
Below is the exact prompt template used to instruct the local LLM to produce the paraphrase of the document:
@@ A.7 Paraphrasing Prompt Template
A.8 Generated paraphrases using various methods on an example document
@@
A.9 Prompt used for LLM as a judge
@@
A.10 Support counts for LLM as a judge
@@
A.11 Full results - DP-Fusion (Multi-Group)
@@
A.12 Full results - DP - Prompt
@@
A.13 Full results - DP - Decoding
@@
A.14 Epsilon vs Attack Success Rates for the Perplexity Attack.
@@
A.15 Epsilon vs Attack Success Rates for the MIN-K Attack.
@@
A.16 Performance of existing Named Entity Recognition Systems
PII tagging is important, as it determines which tokens are covered under the theoretical privacy guarantee. For privacy, recall is the primary concern, while precision mainly impacts utility. Lower precision can, in fact, increase our method’s advantage over the public baseline, as more tokens are treated as private and protected. In the absence of golden labels, we would expect the gap between the public baseline and our method to widen, since higher recall can be achieved in exchange for lower precision. To evaluate the performance of existing taggers, we use the widely adopted Microsoft Presidio library microsoft_presidio_docs, which has been used in prior work staab2024large. We select the best-performing models available within the Presidio suite: BERT-NER (dslim/bert_base_NER), SpaCy (en_core_web_lg), and Flair (flair/ner_english_ontonotes_large).
To test PII-tagging performance on the TAB-ECHR dataset (Section 5.1), we use the same document subset employed in the evaluation of DP-Fusion and focus on identifying PII of type PERSON. We select this entity type because it appears consistently across all considered documents (total $690$ mentions) and includes personal names, which are generally harder to identify than categories like DATES or CODE pham2025can. This is particularly true in the ECHR context, where names tend to be unique, making it difficult for rule-based systems to detect them reliably.

### v4

@@ 4.3 Empirical Privacy Attacks
This section describes empirical attacks to measure lower bounds on the attacker’s advantage measured in the token recovery game, whereas Section˜4.2 provides the theoretical privacy analysis. The token-recovery game states that when attacking a privacy group j{j}, the attacker’s goal is to predict, from the candidate set $C_{j^{*}}$, which (ordered) token set was present in the original input document $D$ which was used as an input to produce the privatized document $D^{\prime}$. Essentially, both attacks aim to model the probability of observing a given paraphrase conditioned on which secret token set was present in the input. This is analogous to prior Membership Inference Attacks (MIA) on LLMs and we can evaluate prior works such as the Min-K Attack (Shi et al., 2023), and the standard baseline LOSS Attack (Yeom et al., 2018), described as follows.
Min-K Attack (Shi et al., 2023). The inference attack, Min-K% (Shi et al., 2023), calculates the average log-likelihood of the k% least-probable tokens i.e., the tokens with the lowest predicted probabilities $\Pr(x_{i}\mid x_{<i})$ in a sequence $x=(x_{1},x_{2},\dots,x_{N})$: $\text{MIN-K\% PROB}(x)=\frac{1}{|\text{Min-K\%}(x)|}\sum_{x_{i}\in\text{Min-K\%}(x)}\log\,\Pr\bigl(x_{i}\,\big|\;x_{1},\dots,x_{i-1}\bigr).$
The LOSS attack (Yeom et al., 2018). This attack calculates a loss for each of the candidate secrets using a surrogate model and the paraphrased document and uses it as a score to predict the true secret.
@@ 5 Experiments
Our experiments use the Qwen 2.5 7B-Instruct model (Qwen et al., 2025) running on a single A100 GPU. This model performs best among our tested set of models (Appendix A.25). We replicate the DPI baseline methods DP-Decoding and DP-Prompt using their publicly released code. To provide a comprehensive overview, we measure utility with multiple metrics: (i) perplexity, computed via teacher forcing on the ground truth document $D$, and (ii) LLM-as-a-judge win-rate, with GPT-4o-mini judge, to compare pairs of generated paraphrases from different methods. We measure privacy through: (i) an upper bound with theoretical guarantees $(\epsilon)$ and (ii) as a lower bound through the adversary’s success rate in the token-recovery game.
@@ 5.1 Experimental Setup
We focus on TAB-ECHR (Pilán et al., 2022) which is a hand-annotated collection of European Court of Human Rights (ECHR) cases (Chalkidis et al., 2019), where private information of eight types (PERSON, CODE, LOC, ORG, DEM, DATETIME, QUANTITY, MISC) is marked. We refer to Section˜A.5 in the Appendix for more details.
@@
2) Empirical Defenses. To simulate simple NER and prompt-engineering baselines (Sec. 2.1), we include two other defenses: No DPI - NER and No DPI - Original Document, where the LLM directly paraphrases the document using only the public tokens $X_{p}$ or the full prompt $D$, respectively (Sec. 4.1). As such approaches will typically involve manually updating the prompt to improve privacy in practical settings, we modify the base prompt (Appendix A.8) with instructions like “produce a natural paraphrase of this for ensuring privacy.” This is only the part of the prompt that is added on top of the existing engineered prompt (described in Appendix A.8) to specifically ask the LLM to ensure privacy. For TAB-ECHR (Sec. 5.1), private tokens are already hand-labeled, so we use these labels directly instead of running an NER system.
@@ 5.3 Utility measured by LLM-as-a-Judge
While perplexity measures token-level fit on the original document $D$, it does not reflect the quality of the generated paraphrase $D^{\prime}$ (Examples in Appendix A.9). Hence, we also evaluate utility with the LLM-as-a-judge setup (Gu et al., 2025) We provide the judge, GPT-4o-mini, with the original document and a pair of paraphrases from different methods or settings, and prompt it (Appendix A.10) to select the paraphrase that retains more information from the original document. We report the resulting win rates in Figure 4, with full support counts for the comparisons in Appendix A.11. For DP-Prompt, we only report the results with $width=50$, as the $width=5$ setting consistently yields garbled outputs, both upon inspection (Appendix A.9) and as indicated by the high perplexity score (Figure 5). DP-Fusion substantially improves over other DPI baselines in this evaluation. Even at the strong privacy (lowest utility setting) ($\alpha\beta=0.01$), it outperforms DP-Decoding and DP-Prompt with $\geq 95\%$ win rate on all settings, except DP-Prompt at $T=0.75$. However, this setting of DP-Prompt is unusable in privacy-focused scenarios, as it provides very low empirical privacy, which we observe in the following section. DP-Fusion, surpasses the public baseline $45\%$ of the time, and at $\alpha\beta=0.1$, it exceeds the public baseline (56% win rate). Within DP-Fusion, stronger privacy ($\alpha\beta=0.01$) yields a lower LLM-as-a-judge win rate than weaker privacy ($\alpha\beta=0.1$), illustrating the expected privacy/utility trade-off. We also evaluate the downstream performance of our generated paraphrases in Appendices A.23 and A.24.
@@ 5.4 Lower Bounds on Privacy
The ASR of the attacks described in Sec. 4.3, together with the corresponding perplexity values of each method (Sec. 5.2), are showcased in Table 1. For each defense method, we report results in two configurations: the highest-utility (lowest-privacy) and the lowest-utility (highest-privacy) settings, as implemented in Section 5.1. We can see that DP-Fusion achieves a $6\times$ higher utility as the best baseline DPI method DP-Prompt at comparable privacy levels (1.426 for $\alpha\beta_{i}=0.10$ versus 8.44 for $\text{w=}50,\text{T=}1.75$). Full results across all parameter settings are presented in Appendices A.12, A.14, and A.13. Additionally, ASR versus $\epsilon$ plots are in Appendix A.15 and A.16.
@@
LOSS-based attack has the highest ASR across all settings. In the strictest privacy setting ($\alpha\beta_{i}=0.01$), DP-Fusion achieves a perplexity (1.459), which is nearly identical to the No DPI - NER baseline (1.46), while maintaining a lower ASR (0.26 vs. 0.2767), thereby offering slightly better utility/privacy tradeoff, but with formal DP guarantees. We believe the slightly better privacy performance of DP-Fusion compared to No-DPI NER arises from the additional randomness introduced during distribution mixing (Algorithm 1), which adds noise and marginally reduces ASR. This effect, lower ASR at the highest-privacy setting ($\alpha\beta_{i}$ = 0.01), also appears in the single-group implementation (Appendix A.19) and on a different dataset (Appendix A.20). However, the difference remains small in all cases.
In the more relaxed setting ($\alpha\beta_{i}=0.10$), DP-Fusion improves utility (PPL = 1.426) with only a marginal increase in ASR (+3.3%). On the other-hand, baseline DPI methods, DP-Decoding and DP-Prompt exhibit significantly higher perplexity (e.g., >100 for DP-Prompt with width 5), indicating heavily degraded outputs. Although DP-Prompt with width 50 and $T=0.75$ achieves lower perplexity (4.26) and produces good quality paraphrases (Figure 4), it does so at the cost of high $\varepsilon$ values (>100,000) and ASR (around 50%), thus providing almost no formal or empirical privacy guarantee.
@@ 6 Discussion
Role of Tagging. DP-Fusion assumes a tagger to define privacy groups; its DP guarantees apply only to the tagged spans. Thus, low false negatives (FN) are required for coverage, not for the validity of the mechanism. We acknowledge that reliance on a fixed PII tagger introduces missed spans, which fall outside the theoretical guarantees; the analysis therefore assumes expert annotation that identifies all PII, even if precision is low. On TAB-ECHR, off-the-shelf taggers (Microsoft, 2025), already achieve low FN ($3.9\%$ FN, $85.4\%$ F1, Appendix A.17) and can be tuned further. With real-world taggers in pipeline, the gap between DP-Fusion and No-DPI NER widens, since we can compensate for tagger imperfections by tuning assigned $\epsilon$ (Appendix A.18). Therefore, developing PII taggers is orthogonal to our work, and DP-Fusion benefits from developments in better NER systems.
@@
Single-Group Implementation. Although we support per-group $\epsilon$, this requires computing m+1m{+}1 distributions per step (1 public + $m$ private), making inference $\approx(m{+}1)\times$ heavier in memory and compute. Increasing $m$ tightens the theoretic privacy (per-group $\epsilon$ decreases with $m$; Thm. 4), but it also increases the effective weight of the public distribution in $p_{\text{final}}$, i.e., more of the public view leaks through. In practice, these effects make the multi-group variant less smooth: as $m$ grows, the fused distribution is increasingly dominated by $p_{\text{pub}}$, so the transition from paraphrasing $D$ to paraphrasing $D\setminus T_{\text{priv}}$ does not vary smoothly with $\epsilon$. We therefore also implement single-group DP-Fusion with one shared $\epsilon$ (Appendix A.19). This variant is more efficient and yields a smoother privacy–utility curve at the expense of weaker theoretical guarantees. We hypothesize that dataset characteristics also contribute. On a different medical PII dataset with more private tokens that impact output paraphrases (Caufield, 2020), we observe smoother privacy–utility trade-offs and a larger gap to No-DPI NER baseline (Appendix A.20).
@@
DP-Fusion against Prompt Injection Attacks. LLMs can be augmented by external databases (e.g., via RAG or web search tools). Given a query, they retrieve multiple chunks from different, potentially untrustworthy sources, which makes the LLM vulnerable to prompt injection attacks (Liu et al., 2024). We use a RAG pipeline and poison a single chunk to jailbreak, achieving an attack success rate (ASR) of $\geq 90\%$ against undefended models. Since the provider knows which chunks came from which source, they can label each chunk as a ’privacy group’ and provably bound the influence of any chunk using DP-Fusion. The defender can control a security/utility trade-off against prompt injection that gracefully degrades toward the no-defense level for larger $\alpha\beta$ values. We evaluate DP-Fusion for different $\alpha\beta$ values and find that at 0.001, 0.01, and 1.0 it provides perfect security ($0\%$ ASR). Full details are in Appendix A.21.
Comparison with baselines. As shown in Table 1, the utility gap between DP-Fusion and the No-DPI NER baseline is modest. This is expected: when few private tokens appear in the source, No-DPI NER removes little content, and DP-Fusion has limited opportunity to preserve additional utility. As the density of sensitive tokens increases, however, DP-Fusion can retain partial information from each private span, whereas No-DPI NER discards all of it. Although constrained by dataset availability, we evaluate a single-group setting (m=1m{=}1) and an alternative dataset (Appendix A.20 and A.20 respectively), where we observe the gap widening.
@@ 7 Conclusion
Ethics Statement
All personally identifiable information (PII) used in this work comes from datasets that were publicly released by their respective owners with the necessary legal clearances and stakeholder consent. We do not collect, annotate, or release any additional PII beyond these existing resources.
Reproducibility Statement
We release a PyPI package alongside a GitHub repository that enables private-span detection and differentially private inference with arbitrary LLMs. The package provides end-to-end support for applying our method to new models and datasets, ensuring full reproducibility of our experiments.
@@ Appendix A Appendix
A.1 LLM Writing Disclosure:
We occasionally used LLMs to paraphrase sentences, proofread text, identify related work and help coding experiments.
A.2 Bisection Search
@@
A.3 Proof of Monotonicity of Rényi Divergence
@@
A.4 Monotonicity of Divergence in $\lambda$
@@
A.5 Detailed information about the TAB-ECHR dataset
@@
  • Quasi identifiers: Publicly known information that doesn’t enable re-identification in isolation but may do so when combined with other quasi-identifiers in the same context. For example, the combination of gender, birth date, and postal code can uniquely identify 63-87% of the U.S. population [Golle, 2006].
@@
A.6 Ablation on candidate set size
@@
A.7 Comparison between data dependent and theoretical $\epsilon$, from $\alpha\beta$
@@
A.8 Paraphrasing Prompt Template
Below is the exact prompt template used to instruct the local LLM to produce the paraphrase of an input document:
@@ A.8 Paraphrasing Prompt Template
A.9 Generated paraphrases using various methods on an example document
@@
A.10 Prompt used for LLM as a judge
@@
A.11 Support counts for LLM as a judge
@@
A.12 Full results - DP-Fusion (Multi-Group)
@@
A.13 Full results - DP - Prompt
@@
A.14 Full results - DP - Decoding
@@
A.15 Epsilon vs Attack Success Rates for the Perplexity Attack.
@@
A.16 Epsilon vs Attack Success Rates for the MIN-K Attack.
@@
A.17 Performance of existing Named Entity Recognition Systems
PII tagging is important, as it determines which tokens are covered under the theoretical privacy guarantee. For privacy, recall is the primary concern, while precision mainly impacts utility. Lower precision can, in fact, increase our method’s advantage over the public baseline, as more tokens are treated as private and protected. In the absence of golden labels, we would expect the gap between the public baseline and our method to widen, since higher recall can be achieved in exchange for lower precision. To evaluate the performance of existing taggers, we use the widely adopted Microsoft Presidio library [Microsoft, 2025], which has been used in prior work [Staab et al., 2024b]. We select the best-performing models available within the Presidio suite: BERT-NER (dslim/bert_base_NER), SpaCy (en_core_web_lg), and Flair (flair/ner_english_ontonotes_large).
To test PII-tagging performance on the TAB-ECHR dataset (Section 5.1), we use the same document subset employed in the evaluation of DP-Fusion and focus on identifying PII of type PERSON. We select this entity type because it appears consistently across all considered documents (total $690$ mentions) and includes personal names, which are generally harder to identify than categories like DATES or CODE [Pham et al., 2025]. This is particularly true in the ECHR context, where names tend to be unique, making it difficult for rule-based systems to detect them reliably.